OpenStudy (anonymous):

Which is not a justification for the proof? A. Pieces of Right Triangles Similarity Theorem B. Side-Side-Side Similarity Theorem C. Substitution D. Addition Property of Equality

OpenStudy (anonymous):

Given: ΔABC is a right triangle. Prove: a2 + b2 = c2 The two-column proof with missing justifications proves the Pythagorean Theorem using similar triangles: Statement Justification Draw an altitude from point C to Line segment AB Let segment BC = a segment CA = b segment AB = c segment CD = h segment DB = x segment AD = y y + x = c c over a equals a over y and c over b equals b over x a2 = cy; b2 = cx a2 + b2 = cy + b2 a2 + b2 = cy + cx a2 + b2 = c(y + x) a2 + b2 = c(c) a2 + b2 = c2

OpenStudy (anonymous):

ignore the jerk. And let me read this proof.

OpenStudy (anonymous):

What do you think it is hun ?

OpenStudy (anonymous):

SSS theorem would be used to prove s triangles are congruent, are we proving that here @Moo_Moo17 ?

OpenStudy (anonymous):

sorry guys I had to reload my internet, its running slow but @liteLegacy i do believe SSS theorem is being used in the trangle

OpenStudy (anonymous):

It okay ^^. and I agree, so we know that is not the answer. let's look at the other answer choices.

OpenStudy (anonymous):

A. Pieces of Right Triangles Similarity Theorem, what do you think of this answer choice ?

OpenStudy (anonymous):

@Moo_Moo17

OpenStudy (anonymous):

it looks like it would be a good answer knowing that the triangle looks likes its a right angle triangle

OpenStudy (anonymous):

That is true but the answer is not right because we "NEVER" use "PIECES" of a theorem, we use the whole theorem or we don't use it at all. :P

OpenStudy (anonymous):

ohhh ok that makes more since now

OpenStudy (anonymous):

Awesome ^^.

OpenStudy (anonymous):

^-^ ok so with tht one crossed out would it be B since i think the triangle is SSS

OpenStudy (anonymous):

@LiteLegacy

OpenStudy (anonymous):

isn't which NOT justification for proof?.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Well were no done let's look at answer D. :P

OpenStudy (anonymous):

ok (: D looks like it would be good what do u think

OpenStudy (anonymous):

@LiteLegacy

OpenStudy (anonymous):

Nah if you look at the two column proof Addition Property of Equality work is done .

OpenStudy (anonymous):

oh i see what mean so is it C then since D was already solved

OpenStudy (anonymous):

C was done in the problem too because B2=cx cx is enter in the place of b2

OpenStudy (anonymous):

oh lol ok so um.. with it be B then ?

OpenStudy (anonymous):

nah A . since we would need sss to prove congruence and stuff :P

OpenStudy (anonymous):

oh ok thank you so much (: do u mind help ing me with a few more??

OpenStudy (anonymous):

Np and sure ^^

OpenStudy (anonymous):

thank you ^w^ here ill post it another question